a) find   b) time at   c) acceleration at   d) velocity just before it reaches ground level...

y=-5t^{3}+15t

a) find  y_{max}

b) time at  y=y_{max}

c) acceleration at  y=y_{max}

d) velocity just before it reaches ground level

e) distance traveled when  y=0

f) time at  y=0 (not  t=0 solution)
Answer #1

Given y -5t3 + 15t

a) To find maximum value of y, we need to set : dt =0

d(-5t3 +15t) dt

15t2 15 = 0

small t = 1s

So at t = 1s, y = ymax = -5(1)3 + 15(1) = 10 m

b) Time at y = ymax = 1s (as found from above)

c) To find acceleration we need:

d2 dt2

d-15t2 +15

dy dt 30t

So acceleration at y = ymax = 301)-30m/s

d) It takes 1 s to reach ymax, so it takes another 1 s to reach the ground level

so total time = 1 + 1 = 2s

velocity just before it reaches ground level = dy dt 25(1545m/

e) Total distance travelled = 2 x ymax = 2 x 10 = 20m

f) time at  y=0 (not  t=0 solution) = 2 s ( as found in (d))